# A Multi-Messenger Search for the Supermassive Black Hole Binary in 3C 66B with the Parkes Pulsar Timing Array

### v1

@@ Abstract
A sub-parsec supermassive black hole binary at the center of the galaxy 3C 66B is a promising candidate for continuous gravitational wave searches with Pulsar Timing Arrays. In this work, we search for such a signal in the third data release of the Parkes Pulsar Timing Array. Matching our priors to estimates of binary parameters from electromagnetic observations, we find log Bayes factor $\ln\mathcal{B}=0.22$, highlighting that the source can be neither confirmed nor ruled out. We place upper limits at $95\%$ credibility on the chirp mass $\mathcal{M}<6.05\times 10^{8}\ M_{\odot}$, and on the characteristic strain amplitude $\textrm{log}_{10}(h_{0})<-14.44$. This partially rules out the parameter space suggested by electromagnetic observations of 3C 66B. We also independently reproduce the calculation of the chirp mass with the 3-mm flux monitor data from the unresolved core of 3C 66B. Based on this, we outline a new methodology for constructing a joint likelihood of electromagnetic and gravitational-wave data from supermassive black hole binaries. Finally, we suggest that firmly established SMBHB candidates may be treated as standard sirens for complementary constraints on the universe expansion rate. With this, we obtain constraints on the Hubble constant with 3C 66B.
@@
software: ptmcmcsampler (Ellis & van Haasteren, 2019), enterprise (Ellis et al., 2019), enterprise_warp at github.com/bvgoncharov/enterprise_warp, enterprise_extensions (Taylor et al., 2021).
1 Introduction
Pulsar Timing Arrays (PTAs) are experiments for long-term monitoring of millisecond pulsar pulse arrival times with the primary goal of detecting nanohertz-frequency gravitational waves. Recent PTA data show evidence for the nanohertz-frequency gravitational wave background (GWB, Goncharov et al., 2021a; Arzoumanian et al., 2020a; Chen et al., 2021; Antoniadis et al., 2022; Agazie et al., 2023a; EPTA Collaboration et al., 2023; Reardon et al., 2023). The background origin is consistent with the superposition of adiabatically inspiraling supermassive black hole binaries (SMBHBs) (Goncharov et al., 2024). Provided confident detection of the background, the detection of a continuous gravitational wave (CGW) from an individual SMBHB is the next milestone for PTAs. Such an observation would decisively resolve the final parsec problem, which represents difficulties in modeling the evolution of SMBHBs to sub-parsec separations. When the host galaxy of CGW is known, the detection unlocks multi-messenger astronomy and cosmology with PTAs.
One of the most promising sources of CGWs is the supermassive black hole binary candidate in the center of galaxy 3C 66B. By using long-baseline interferometry, Sudou et al. (2003) found evidence for elliptical binary orbital motion in this radio galaxy’s core and inferred a gravitational wave (GW) frequency of $f_{\textrm{GW}}=60.4\pm 1.73$ nHz. Further observations by Iguchi et al. (2010) found that the binary has a chirp mass $\mathcal{M}=7.9^{+3.8}_{-4.5}\times 10^{8}\ M_{\odot}$. Galaxy 3C 66B is located at a redshift $z=0.02126$ (van den Bosch et al., 2015), which corresponds to the luminosity distance $D_{\text{L}}=94.17^{+0.20}_{-0.20}$ Mpc (Planck Collaboration et al., 2020). Based on the above, the expected strain amplitude $h_{0}$ of a CGW at 3C 66B is $7.3^{+6.8}_{-5.8}\times 10^{-15}$. This is the highest known strain amplitude of CGW sources in the PTA frequency band, and is within reach for PTA experiments. The summary of the expected properties of the SMBHB at 3C 66B is shown in Table 1. Because the properties in the table are inferred from electromagnetic (EM) observations, we will also refer to this model of the SMBHB at 3C 66B as the EM model.
Table 1: Summary of properties of the SMBHB candidate at 3C 66B.
@@ 1 Introduction
$\mathcal{M}$ $7.9^{+3.8}_{-4.5}\times 10^{8}\ M_{\odot}$ Iguchi et al. (2010)
$f_{\textrm{GW}}$ $60.4\pm 1.73$ nHz Sudou et al. (2003)
$D_{\text{L}}$ $94.17\pm 0.20$ Mpc van den Bosch et al. (2015)
R.A. 02h 23m 11.4110s Hunt et al. (2021)
Decl. +42d 59m 31.385s Hunt et al. (2021)
@@
It is worth noting that Iguchi et al. (2010) revised the chirp mass of the SMBHB in 3C 66B inferred in the earlier work of Sudou et al. (2003). The initial model of the SMBHB at the center of 3C 66B proposed by Sudou et al. (2003) suggested a very high chirp mass of $1.3\times 10^{10}\ M_{\odot}$. Jenet et al. (2004) has ruled out this model based on seven years of timing from PSR B1855+09 (Kaspi et al., 1994). Assuming a circular SMBHB at 3C 66B, they placed an upper limit of $7\times 10^{9}\ M_{\odot}$ on its chirp mass. This upper limit does not exclude the model by Iguchi et al. (2010), which motivates further studies.
@@
Following the revision of the chirp mass of the SMBHB candidate by Iguchi et al. (2010), a number of searches for a CGW from 3C 66B with PTAs have been performed. The authors of Arzoumanian et al. (2020b) performed a search for an SMBHB in a circular orbit with the 11-year dataset of the North American Nanohertz Observatory for Gravitational Waves (NANOGrav). An upper limit on the chirp mass of $1.65\times 10^{9}\ M_{\odot}$ was placed. A subsequent search was then performed for both circular and eccentric SMBHBs using the NANOGrav 12.5-year dataset. The search for a circular binary (Arzoumanian et al., 2023) constrained the mass to $1.41\times 10^{9}\ M_{\odot}$. The search, which included eccentricity in the SMBHB system (Agazie et al., 2024), placed their tightest constraints on the upper limit of chirp mass as $1.81\times 10^{9}\ M_{\odot}$. The results from the 12.5-year NANOGrav dataset are all still compatible with the Iguchi et al. (2010) model, yet claim no detection, which motivates further searches for the source in pulsar timing array datasets. With NANOGrav 15-year data, the limit on the chirp mass at $95\%$ credibility is $9.54\times 10^{8}~M_{\odot}$ (Agarwal et al., 2025).
@@
In this work, we perform a search for CGW from an SMBHB in a circular orbit at 3C 66B using the third data release (DR3) of the Parkes Pulsar Timing Array (PPTA) (Zic et al., 2023). It is the first targeted search for 3C 66B with the PPTA data. We expect the search based on this data to provide better constraints on CGW from 3C 66B, thanks to an increased time span exceeding $18$ years in $13$ pulsars. The targeting is performed by setting the sky position as a fixed parameter, as well as setting priors for other parameters based on electromagnetic observations summarized in Table 1. We employ both Bayesian and frequentist methods to calculate the significance of the hypothesis that an SMBHB with sub-parsec separation resides in 3C 66B. Finally, we illustrate current capabilities of cosmology with PTAs by constraining the Hubble constant based on the premise that an SMBHB exists at 3C 66B.
@@
The rest of the paper is organized as follows. In Section 2, we introduce the dataset from the third data release of the PPTA. In Section 3, we introduce the signal model, the Bayesian analysis methodology, and our noise models. In Section 4, we present our results and discuss the model selection, upper limits on the chirp mass, upper limits on the characteristic strain amplitude, upper limits as a function of CGW frequency, the frequentist analysis results, and how targeted searches in PTAs can be used for cosmology experiments. In Section 5 we present our conclusions.
2 Data
@@
The Parkes Pulsar Timing Array dataset (Zic et al., 2023) is based on monitoring of $N_{\text{psr}}=32$ millisecond radio pulsars. The total observing span of PPTA DR3 is $T_{\text{obs}}=18$ years. The first observations took place on February 6, 2004, and the last observations took place on March 8, 2022. The observations were conducted using the 64-m Parkes ‘Murriyang’ radio telescope in Australia, and observations were taken for each pulsar approximately every three weeks. The most recent $3$ years of observations in DR3 are carried out using the ultra-wide bandwidth, low-frequency receiver (UWL, Hobbs et al., 2020). In this work, only the data from 31 pulsars are used. PSR J1741+1351 is excluded from the analysis due to the fact that it was only added to the PPTA observations after the commissioning of the UWL and was only observed with low priority. With only 16 observations, this pulsar is not sensitive to any CGW signal and will therefore not contribute to the results.
@@
As with any PTA data, PPTA DR3 is a set of pulse arrival time observations, including observation details such as radio frequency and backend-receiver information, as well as the pulsar ephemeris, known as the timing model. A timing model predicts pulse arrival times based on the pulsar-specific spin, astrometry, properties of the interstellar plasma along the line of sight, and the binary parameters of the pulsar when applicable. We use the DE440 ephemeris (Park et al., 2021) to obtain pulse arrival times in an inertial frame centered at the Solar System Barycenter (SSB).
@@
Additionally, we follow Iguchi et al. (2010) by employing the same millimeter-wavelength flux variation data for 3C 66B. There are two components of this data. The first component of the data is collected at $93.7$ GHz using the Nobeyama Millimeter Array (NoMA) of the National Astronomical Observatory of Japan (NAOJ). The second component of the data is collected at $86.2$ GHz using the Plateau de Bure Interferometer (PdBI) of the Institut de Radioastronomie Millimétrique (IRAM).
3 Methodology
3.1 Signal Model
@@
We use the CGW signal model for non-spinning SMBHBs in circular [^1] Iguchi et al. (2010) point out that 3C 66B may be in a circular orbit. orbits at zeroth post-Newtonian order described in (Aggarwal et al., 2019; Arzoumanian et al., 2020b, 2023). More details of the model are provided in Appendix A. The CGW model for metric perturbations at SSB (the Earth term) is based on $8$ parameters:
$\{\theta,\phi,f_{\text{GW}},\Phi_{0},\psi,i,\mathcal{M},(D_{\text{L}}\wedge h_{0})\}$ (1)
@@
Where $(\theta,\phi)$ is the sky position, $f_{\text{GW}}$ is the gravitational wave frequency, $\Phi_{0}$ is the orbital phase, $\psi$ is the gravitational wave polarization angle, $i$ is the binary inclination angle, $\mathcal{M}$ is the chirp mass, $D_{\text{L}}$ is the luminosity distance and $h_{0}$ is the gravitational strain amplitude. At zeroth post-Newtonian order, we compute the amplitude as
@@
unless it is provided as a free parameter (hence, $D_{\text{L}}\wedge h_{0}$). The full CGW model is based on metric perturbations at both the SSB and each pulsar (pulsar terms). This requires additional $2N_{\text{psr}}$ parameters $\{L_{i},\Phi_{i}\}$, $i=(1,N_{\text{psr}})$. They are the pulsar distance and the orbital phase at the pulsar, respectively. We provide more information on pulsar distance in Appendix C. Searching only for the Earth term is sufficient for detection and upper limits, whereas including the pulsar term enhances the search sensitivity and remains important for parameter estimation when the signal is detected.
We model the evolution of CGW frequency on the timescales of light travel time between pulsars and Earth. However, for simplicity, we neglect frequency evolution on the timescale of PPTA DR3. This approach is referred to as the stationary phase approximation. To demonstrate the validity of this approach, we employ the equation for the evolution of CGW frequency as a function of time at zeroth post-Newtonian order from (Abbott et al., 2017a),
@@
where $f_{\text{GW}}$ is the gravitational wave frequency, $G$ is Newton’s constant, $c$ is the speed of light, and $t_{\text{c}}$ is the time of coalescence. We present the expected frequency evolution of 3C 66B at SSB in Figure 1. Based on the EM model for 3C 66B, a circular SMBHB is only expected to evolve by about $0.04$ nHz/year. In the 20 years since the initial frequency was determined by Sudou et al. (2003), the frequency may have increased by $0.86$ nHz. This width does not exceed the frequency resolution of PPTA DR3 of $T^{-1}_{\text{obs}}=1.75$ nHz. Therefore, according to the model provided by Sudou et al. (2003), the stationary phase approximation is sufficient for our analysis.
Figure 1: The evolution of the CGW frequency of 3C 66B at the SSB and the associated uncertainty from Sudou et al. (2003) is shown by the pink diagonal line. The dashed red line and the shaded red region show the frequency inferred by and their measurement uncertainty, respectively. The vertical red band represents the observing time frame of Sudou et al. (2003). The blue shaded region represents the observing time frame of Iguchi et al. (2010). The green shaded region represents the observing time frame of the PPTA DR3 dataset. The gray shaded region represents one frequency bin in PPTA DR3 based on $T^{-1}_{\text{obs}}$.
3.2 Bayesian inference
We perform parameter estimation and model selection by computing posteriors, $\mathcal{P}(\bm{\theta}|\bm{\delta t})=\mathcal{Z}^{-1}\mathcal{L}(\bm{\delta t}|\bm{\theta})\pi(\bm{\theta})$, where $\mathcal{L}(\bm{\delta t}|\bm{\theta})$ is the likelihood of data time series vector $\bm{\delta t}$ given a vector of model parameters $\theta$ for signals and noise, and $\pi(\bm{\theta})$ is the prior. The constant $\mathcal{Z}$ is the integral of the numerator over $\bm{\theta}$. It is also referred to as the evidence or the fully marginalized likelihood. We neglect $\mathcal{Z}$ in our parameter estimation and obtain posterior samples using parallel-tempered Markov-Chain Monte-Carlo (Ellis & van Haasteren, 2019). The model selection follows from computing Bayes factors, $\mathcal{B}=\mathcal{Z}/\mathcal{Z}_{\varnothing}$, where $\mathcal{Z}_{\varnothing}$ is the evidence of the noise model and $\mathcal{Z}$ is the evidence of the signal model, which includes both the noise and the CGW. Instead of computing $\mathcal{Z}$ separately for the two models, we compute the Bayes factor directly using the product-space sampling method (Hee et al., 2016).

### v2

# A Multimessenger Search for the Supermassive Black Hole Binary in 3C 66B with the Parkes Pulsar Timing Array

@@ Abstract
A subparsec supermassive black hole binary (SMBHB) at the center of the galaxy 3C 66B is a promising candidate for continuous gravitational-wave searches with pulsar timing arrays (PTAs). In this work, we search for such a signal in the third data release of the Parkes Pulsar Timing Array. Matching our priors to estimates of binary parameters from electromagnetic observations, we find a log Bayes factor $\ln\mathcal{B}=-0.0027(7)$, highlighting that the source can be neither confirmed nor ruled out. We place upper limits at $95\%$ credibility on the chirp mass $\mathcal{M}<6.90\times 10^{8}\ M_{\odot}$, and on the characteristic strain amplitude $\textrm{log}_{10}(h_{0})<-14.44$. This partially rules out the parameter space suggested by electromagnetic (EM) observations of 3C 66B. We also independently reproduce the calculation of the chirp mass with the 3 mm flux monitor data from the unresolved core of 3C 66B. Based on this, we outline a new methodology for constructing a joint likelihood of EM and gravitational-wave data from SMBHBs. Finally, we suggest that targeted searches may allow firmly established SMBHB candidates to be treated as standard sirens, for complementary constraints on the Universe expansion rate.
@@
software: ptmcmcsampler (Ellis and van Haasteren, 2019), enterprise (Ellis et al., 2019), enterprise_warp at github.com/bvgoncharov/enterprise_warp, enterprise_extensions (Taylor et al., 2021).
I Introduction
Pulsar timing arrays (PTAs) are experiments for long-term monitoring of millisecond pulsar pulse arrival times with the primary goal of detecting nanohertz-frequency gravitational waves (GWs). Recent PTA data show evidence for a nanohertz-frequency gravitational wave background (GWB; Goncharov et al., 2021b; Arzoumanian et al., 2020a; Chen et al., 2021; Antoniadis et al., 2022; Agazie et al., 2023a; EPTA Collaboration et al., 2023; Reardon et al., 2023). The background origin is consistent with the superposition of adiabatically inspiraling supermassive black hole binaries  (SMBHBs; Goncharov et al., 2025). Provided confident detection of the background, the detection of a continuous gravitational wave (CGW) from an individual SMBHB is the next milestone for PTAs. Such an observation would decisively resolve the final parsec problem, which represents difficulties in modeling the evolution of SMBHBs to subparsec separations (Milosavljević and Merritt, 2003). When the host galaxy of CGW source is known, the detection unlocks multimessenger astronomy and cosmology with PTAs (Schutz, 1986; Holz and Hughes, 2005; Charisi et al., 2022).
One of the most promising sources of CGWs is the SMBHB candidate in the center of galaxy 3C 66B. By using long baseline interferometry, Sudou et al. (2003) found evidence for elliptical binary orbital motion in this radio galaxy’s core and inferred a GW frequency of $f_{\textrm{GW}}=60.4\pm 1.73$ nHz. Further observations by Iguchi et al. (2010) found that the binary has a chirp mass $\mathcal{M}=7.9^{+3.8}_{-4.5}\times 10^{8}\ M_{\odot}$, primarily through measuring the orbital radius of the electromagnetic (EM)-emitting black hole and the semi-major axis of an SMBHB. Galaxy 3C 66B is located at a redshift $z=0.02126$ (van den Bosch et al., 2015), which corresponds to a luminosity distance $D_{\text{L}}=94.17^{+0.20}_{-0.20}$ Mpc (Planck Collaboration et al., 2020).
Based on the above, the expected strain amplitude $h_{0}$ of a CGW from 3C 66B is $7.3^{+6.8}_{-5.8}\times 10^{-15}$. This is the highest known strain amplitude of CGW sources in the PTA frequency band, and is within reach for PTA experiments. A summary of the expected properties of the SMBHB in 3C 66B is shown in Table 1. Because the properties in the table are inferred from EM observations, we will also refer to this model of the SMBHB in 3C 66B as the “EM model”.
Table 1: Summary of the Properties of the Supermassive Black Hole Binary Candidate in 3C 66B.
@@ I Introduction
$\mathcal{M}$ ($M_{\odot}$) $7.9^{+3.8}_{-4.5}\times 10^{8}$ Iguchi et al. (2010)
$f_{\textrm{GW}}$ (nHz) $60.4\pm 1.73$ Sudou et al. (2003)
$D_{\text{L}}$ (Mpc) $94.17\pm 0.20$ van den Bosch et al. (2015)
R.A. (HH:MM: SS.SSS) 02:23:11.4110 Hunt et al. (2021)
Decl. (DD:MM: SS.SSS) +42:59:31.385 Hunt et al. (2021)
@@
It is worth noting that Iguchi et al. (2010) revised the chirp mass of the SMBHB in 3C 66B inferred in the earlier work of Sudou et al. (2003). The initial model of the SMBHB at the center of 3C 66B proposed by Sudou et al. (2003) suggested a very high chirp mass of $1.3\times 10^{10}\ M_{\odot}$. Jenet et al. (2004) have ruled out this model based on 7 yr of timing from PSR B1855+09 (Kaspi et al., 1994). Assuming a circular SMBHB in 3C 66B, they placed an upper limit of $7\times 10^{9}\ M_{\odot}$ on its chirp mass. This upper limit does not exclude the model by Iguchi et al. (2010), which motivates further studies.
@@
Following the revision of the chirp mass of the SMBHB candidate by Iguchi et al. (2010), a number of searches for a CGW from 3C 66B with PTAs have been performed. The authors of Arzoumanian et al. (2020b) performed a search for an SMBHB in a circular orbit with the 11 yr dataset of the North American Nanohertz Observatory for Gravitational Waves (NANOGrav). An upper limit on the chirp mass of $1.65\times 10^{9}\ M_{\odot}$ was placed. A subsequent search was then performed for both circular and eccentric SMBHBs using the NANOGrav 12.5 yr dataset. The search for a circular binary (Arzoumanian et al., 2023) constrained the mass to $1.41\times 10^{9}\ M_{\odot}$. The search, which included eccentricity in the SMBHB system (Agazie et al., 2024), placed their tightest constraints on the upper limit of the chirp mass with the value of $1.81\times 10^{9}\ M_{\odot}$. The results from the 12.5 yr NANOGrav dataset are all still compatible with the Iguchi et al. (2010) model, yet claim no detection, which motivates further searches for the source in PTA datasets. Similar to the results shown later in our work, the NANOGrav 15 yr dataset places a limit on the chirp mass at $95\%$ credibility as $9.54\times 10^{8}~M_{\odot}$ (Agarwal et al., 2025). On this same dataset, Tian et al. (2025) found that for an eccentric model, the upper limit on the chirp mass is of $1.82\times 10^{9}~M_{\odot}$.
@@
In this work, we perform a search for a CGW from an SMBHB with a circular orbit in 3C 66B using the third data release (DR3) of the Parkes Pulsar Timing Array  (PPTA; Zic et al., 2023). It is the first targeted search for 3C 66B with PPTA data. We expect the search based on these data to provide better constraints on a CGW from 3C 66B, thanks to an increased time span exceeding 18 yr in 13 pulsars (Mingarelli et al., 2017). As motivated by Liu and Vigeland (2021), the targeting is performed by setting the sky position as a fixed parameter, as well as setting priors for other parameters based on the EM observations summarized in Table 1. We employ both Bayesian and frequentist methods to calculate the significance of the hypothesis that an SMBHB with subparsec separation resides in 3C 66B. Finally, we discuss the current feasibility of cosmology with PTAs and produce constraints on the Hubble–Lemaître constant, assuming that the EM model is valid.
@@
The rest of the paper is organized as follows. In Section II, we introduce the dataset from DR3 of the PPTA. In Section III, we introduce the signal model, the Bayesian analysis methodology, and our noise models. In Section IV, we present our results and discuss the model selection, upper limits on the chirp mass, upper limits on the characteristic strain amplitude, upper limits as a function of CGW frequency, and the frequentist analysis results. In Section V, we discuss new methods for a joint-likelihood analysis of 3C 66B and whether targeted searches in PTAs can be used for cosmology experiments. In Section VI, we present our conclusions.
II Data
@@
The PPTA DR3 dataset (Zic et al., 2023) is based on monitoring of $N_{\text{psr}}=32$ millisecond radio pulsars. The total observing span of PPTA DR3 is $T_{\text{obs}}=18$ yr. The first observations took place on 2004 February 6, and the last observations took place on 2022 March 8. The observations were conducted using the 64 m Parkes ‘Murriyang’ radio telescope in Australia, and observations were taken for each pulsar approximately every 3 weeks. The most recent 3 yr of observations in DR3 are carried out using the ultra-wide bandwidth, low-frequency receiver (UWL; Hobbs et al., 2020). In this work, only the data from 31 pulsars are used. PSR J1741+1351 is excluded from the analysis due to the fact that it was only added to the PPTA observations after the commissioning of the UWL and was only observed with low priority. With only 16 observations, this pulsar is not sensitive to any CGW signal and will therefore not contribute to the results.
@@
As with any PTA data, PPTA DR3 is a set of pulse arrival time observations, including observation details such as radio frequency and backend-receiver information, as well as the pulsar ephemeris, known as the timing model. A timing model predicts pulse arrival times based on the pulsar’s specific spin, astrometry, properties of the interstellar plasma along the line of sight, and the binary parameters of the pulsar when applicable. We use the DE440 ephemeris (Park et al., 2021) to obtain pulse arrival times in an inertial frame centered at the solar system barycenter (SSB).
@@
Additionally, we follow Iguchi et al. (2010) by employing the same millimeter-wavelength flux variation data for 3C 66B. There are two components of this data. The first component of the data is collected at $93.7$ GHz using the Nobeyama Millimeter Array of the National Astronomical Observatory of Japan. The second component of the data is collected at $86.2$ GHz using the Plateau de Bure Interferometer of the Institut de Radioastronomie Millimétrique.
III Methodology
III.1 Signal Model
@@
We use the CGW signal model for nonspinning SMBHBs in circular [^1] Iguchi et al. (2010) point out that 3C 66B may be in a circular orbit. orbits at zeroth post-Newtonian order described in Aggarwal et al. (2019) and Arzoumanian et al. (2020b, 2023). More details of the model are provided in Appendix A. The CGW model for metric perturbations at SSB (the Earth term) is based on eight parameters:
$\{\theta,\phi,f_{\text{GW}},\Phi_{0},\psi,\iota,\mathcal{M},(D_{\text{L}}\wedge h_{0})\}$ (1)
@@
where $(\theta,\phi)$ is the sky position, $f_{\text{GW}}$ is the GW frequency, $\Phi_{0}$ is the orbital phase, $\psi$ is the GW polarization angle, $\iota$ is the binary inclination angle, $\mathcal{M}$ is the chirp mass, $D_{\text{L}}$ is the luminosity distance and $h_{0}$ is the gravitational strain amplitude. At zeroth post-Newtonian order, we compute the amplitude as
@@
unless it is provided as a free parameter (hence, $D_{\text{L}}\wedge h_{0}$). The full CGW model is based on metric perturbations at both the SSB and each pulsar (pulsar terms). This requires additional $2N_{\text{psr}}$ parameters $\{L_{i},\Phi_{i}\}$, $i=(1,N_{\text{psr}})$. They are the pulsar distance and the orbital phase at the pulsar, respectively. We provide more information on the pulsar distances in Appendix C. Searching only for the Earth term is sufficient for detection and upper limits, whereas including the pulsar term enhances the search sensitivity and remains important for parameter estimation when the signal is detected.
We model the evolution of CGW frequency on the timescales of the light travel time between the pulsars and Earth. However, for simplicity, we neglect frequency evolution on the timescale of PPTA DR3. This approach is referred to as the stationary phase approximation. To demonstrate the validity of this approach, we employ an equation for the evolution of CGW frequency as a function of time at zeroth post-Newtonian order from (Abbott et al., 2017a):
@@
where $f_{\text{GW}}$ is the GW frequency, $G$ is Newton’s constant, $c$ is the speed of light, and $t_{\text{c}}$ is the time of coalescence. We present the expected frequency evolution of 3C 66B at SSB in Figure 1. Based on the EM model for 3C 66B, a circular SMBHB is only expected to evolve by about $0.04$ $\text{nHz}\ \text{yr}^{-1}$. In the 20 yr since the initial frequency was determined by Sudou et al. (2003), the frequency may have increased by $0.86$ nHz. This width does not exceed the frequency resolution of PPTA DR3 of $T^{-1}_{\text{obs}}=1.75$ nHz. Therefore, according to the model provided by Sudou et al. (2003), the stationary phase approximation is sufficient for our analysis.
Figure 1: The evolution of the CGW frequency of 3C 66B at the SSB, and the associated uncertainty from Sudou et al. (2003) is shown by the pink diagonal line. The dashed red line and the shaded red region show inferred frequency and their measurement uncertainty, respectively. The vertical red band represents the observing time frame of Sudou et al. (2003). The blue shaded region represents the observing time frame of Iguchi et al. (2010). The green shaded region represents the observing time frame of the PPTA DR3 dataset. The gray shaded region represents one frequency bin in PPTA DR3 based on $T^{-1}_{\text{obs}}$.
III.2 Bayesian Inference
We perform parameter estimation and model selection by computing posteriors, $\mathcal{P}(\bm{\theta}|\bm{\delta t})=\mathcal{Z}^{-1}\mathcal{L}(\bm{\delta t}|\bm{\theta})\pi(\bm{\theta})$, where $\mathcal{L}(\bm{\delta t}|\bm{\theta})$ is the likelihood of data time series vector $\bm{\delta t}$ given a vector of model parameters $\theta$ for signals and noise, and $\pi(\bm{\theta})$ is the prior. The constant $\mathcal{Z}$ is the integral of the numerator over $\bm{\theta}$. It is also referred to as the evidence or the fully marginalized likelihood. We neglect $\mathcal{Z}$ in our parameter estimation and obtain posterior samples using parallel-tempered Markov-Chain Monte-Carlo (Ellis and van Haasteren, 2019). The model selection follows from computing the Bayes factor, $\mathcal{B}=\mathcal{Z}/\mathcal{Z}_{\varnothing}$, where $\mathcal{Z}_{\varnothing}$ is the evidence of the noise model and $\mathcal{Z}$ is the evidence of the signal model, which includes both the noise and the CGW. Instead of computing $\mathcal{Z}$ separately for the two models, we compute the Bayes factor directly using the product-space sampling method (Hee et al., 2016).